# High-dimensional sliced inverse regression with endogeneity

### v1

@@ Abstract
Sliced inverse regression (SIR) is a popular sufficient dimension reduction method that identifies a few linear transformations of the covariates without losing regression information with the response. In high-dimensional settings, SIR can be combined with sparsity penalties to achieve sufficient dimension reduction and variable selection simultaneously. Nevertheless, both classical and sparse estimators assume the covariates are exogenous. However, endogeneity can arise in a variety of situations, such as when variables are omitted or are measured with error. In this article, we show such endogeneity invalidates SIR estimators, leading to inconsistent estimation of the true central subspace. To address this challenge, we propose a two-stage Lasso SIR estimator, which first constructs a sparse high-dimensional instrumental variables model to obtain fitted values of the covariates spanned by the instruments, and then applies SIR augmented with a Lasso penalty on these fitted values. We establish theoretical bounds for the estimation and selection consistency of the true central subspace for the proposed estimators, allowing the number of covariates and instruments to grow exponentially with the sample size. Simulation studies and applications to two real-world datasets in nutrition and genetics illustrate the superior empirical performance of the two-stage Lasso SIR estimator compared with existing methods that disregard endogeneity and/or nonlinearity in the outcome model.
@@ 1 Introduction
Sufficient dimension reduction (SDR, Ma and Zhu,, 2013; Li,, 2018) refers to a class of methods that assume the outcome $y\in\mathbb{R}$ depends on a set of covariates $\mathbf{X}\in\mathbb{R}^{p}$ through a small number of linear combinations $\bm{\beta}_{1}^{\top}\mathbf{X},\ldots,\bm{\beta}_{d}^{\top}\mathbf{X}$, with $d\ll p$. These linear combinations are known as sufficient predictors and retain the full regression information between the response and covariates. Since the pioneering work ofLi, (1991), a vast literature has grown on different SDR approaches, ranging from inverse-moment-based and regression-based methods (e.g., Cook and Forzani,, 2008; Bura et al.,, 2022), forward regression (e.g., Xia et al.,, 2002; Dai et al.,, 2024), and semiparametric techniques (e.g., Zhou and Zhu,, 2016; Sheng and Yin,, 2016). Among the various SDR methods, sliced inverse regression (SIR, Li,, 1991) remains the most widely used due to its simplicity and computational efficiency, and as such continues to be the subject of much research (see Tan et al.,, 2018; Xu et al.,, 2022, among many others). When the number of covariates $p$ is large, it is often (further) assumed each linear combination depends only on a few covariates, i.e., the vectors $\bm{\beta}_{1},\ldots,\bm{\beta}_{d}$ are sparse. Recent literature has thus sought to combine SIR and other SDR methods with variable selection to achieve sparse sufficient dimension reduction (e.g., Lin et al.,, 2018, 2019; Nghiem et al., 2023c, ).
Much of the literature dedicated to SIR and many other SDR methods, including those mentioned above, has focused on the setting where all covariates are exogenous, and little has been done to tackle sufficient dimension reduction with endogeneity. While there is some related literature on single index models with endogenity (Hu et al.,, 2015; Gao et al.,, 2020), their approach requires approximation of the link function and is not easy to generalize to multiple index models. More specifically, letting $\mathbf{B}=[\bm{\beta}_{1},\ldots,\bm{\beta}_{d}]$ denote a $p\times d$ matrix, we can formulate SDR as a multiple index model
@@
where $f$ denotes an unknown link function, and $\varepsilon\in\mathbb{R}$ is random noise. In (1), the target of estimation is the column space of $\mathbf{B}$, also known as the central subspace. Almost all current SDR methods assume the random noise $\varepsilon$ is independent of $\mathbf{X}$, implying $y\perp\mathbf{X}\mid\mathbf{B}^{\top}\mathbf{X}$ where $\perp$’ denotes statistical independence. However, this assumption may not be reasonable in many settings where the random noise $\varepsilon$ is dependent upon $\mathbf{X}$ i.e., the covariates are endogenous.
We discuss two examples, corresponding to two data applications in Section 6, where such endogeneity can arise.
  • Omitted-variable bias occurs when one or more (truly important) covariate is not included in the model. A classic example relates to studying the effect of education on future earnings. Here, it is well-known future earnings depend on many other variables that are also highly correlated with education level, such as family background or an individual’s overall abilities (Card,, 1999). Such variables are usually not fully known or not collected, so the ordinary least squares fit of a regression model of earnings on education levels and other observed factors produces a biased estimate for the true effect of education (Card,, 1999). Similarly, in Section 6.1, we consider a genomic study that relates a biological outcome (e.g., body weight) to gene expressions of mice. Outside of gene expressions, there can be potentially unmeasured confounding factors, such as those related to environmental conditions, which influence both gene expression levels and the outcome.
@@
  • Covariate measurement error arises when some or all covariates are not accurately measured. In the second data application in Section 6.2, we consider a nutritional study relating an individual’s health outcomes to their long–term diet, collected through 24–hour recall interviews that are subject to measurement errors (Carroll et al.,, 2006). Let $\mathbf{X}$ denote the long-term dietary covariates. The observed data can then be modeled by $\mathbf{W}=\mathbf{X}+\mathbf{E}$, with $\mathbf{E}$ denoting an additive measurement error independent of $\mathbf{X}$. In this situation, even when in the true model (1) it holds that $\varepsilon\perp\mathbf{X}$, a naive analysis which replaces $\mathbf{X}$ by $\mathbf{W}$ results in endogeneity, since we can write $y=f(\mathbf{B}^{\top}\mathbf{W}+\mathbf{B}^{\top}\mathbf{E},\varepsilon)$ and $\mathbf{E}$ is correlated with $\mathbf{W}$. This causes naive estimators that ignore measurement errors to be biased and inconsistent.
As an aside, note that while the omitted variables lead to endogeneity because too few covariates are included in the model, Fan and Liao, (2014) argued that endogeneity also arises when a large number of covariates are included in the model i.e., $\mathbf{X}$ is high-dimensional in (1). Specifically, for large $p$ there is a high probability at least one covariate is correlated with the random noise. For instance, genome-wide association studies typically include thousands of gene expression covariates to model a clinical outcome, and at least one gene expression is likely to correlate with random noise reflecting experimental conditions or environmental perturbations (Lin et al.,, 2015).
@@
In this paper, we show under fairly general conditions that endogeneity also invalidates sliced inverse regression, leading to inconsistent estimates of the true column space of $\mathbf{B}$. To address this issue, we propose a two-stage Lasso SIR method to estimate the central space in the presence of endogeneity when valid instrumental variables are available. In the first stage, we construct a sparse multivariate linear model regressing all the covariates against the instruments and then obtain fitted values from this model. In the second stage, we apply a sparse sliced inverse regression between the response and the fitted values from the first stage, to perform simultaneous SDR and variable selection.
@@
Addressing endogeneity via instrumental variables has been extensively studied in the econometrics and statistics literature; see Angrist and Krueger, (2001) for a review. Classical instrumental variable estimators are mostly studied with a linear model for the outcome, i.e., $y=\mathbf{X}^{\top}\bm{\beta}_{0}+\varepsilon$, and often computing a two-stage least squares estimator or a limited information maximum likelihood estimator. Hansen et al., (2008) showed in this setting that these estimators are consistent for $\bm{\beta}_{0}$ when the number of instruments $q$ increases slowly with the sample size. More recently however, a number of methods have combined instrumental variables with regularization techniques on the covariates and/or instrumental variables to improve their performance in high-dimensional settings e.g., the Lasso-type generalized methods of moments (Caner,, 2009; Caner et al.,, 2018), focused generalized methods of moments (Fan and Liao,, 2014), and two-stage penalized least squares Lin et al., (2015); the last of these methods inspires this article. Another line of research has focused on the use of sparsity-inducing penalties to estimate optimal instruments from a large set of a-priori valid instruments (see for instance Belloni et al.,, 2012; Chernozhukov et al.,, 2015), or to screen out invalid instruments (e.g., Windmeijer et al.,, 2019). To the best of our knowledge though, none of these methods allow for the presence of non-linear link functions in the model for the response e.g., the multiple index model (1), and more broadly have not been adapted to the SDR context.
@@
Under a setting where both the number of covariates $p$ and the number of instrumental variables $q$ can grow exponentially with sample size $n$, we derive theoretical bounds on the estimation and selection consistency of the two-stage Lasso SIR estimator. Simulation studies demonstrate the proposed estimator exhibits superior finite sample performance compared with a single stage Lasso SIR estimator when endogeneity is present but ignored, and the two-stage regularized estimator proposed by Lin et al., (2015) when the outcome model is non-linear in nature. We apply the proposed method to two real data applications studying the effects of different nutrients on the cholesterol level (covariate measurement error), and for identifying genes that are relevant to the body weight of mice (omitted variable bias).
The remainder of this article is structured as follows. In Section 2, we briefly review two versions of sliced inverse regression and prove that endogeneity causes SIR to be inconsistent for estimating the target subspace. In Section 3, we formulate the proposed two-stage Lasso estimators, and derive their theoretical properties in Section 4. Section 5 presents simulation studies to compare the performance of the proposed estimator with existing methods, while Section 6 discusses two data applications Finally, Section 7 offers some concluding remarks.
@@ 2 SIR and endogeneity
Before proceeding, we set out some notation to be used throughout the paper. For any matrix $\mathbf{A}$ with elements $a_{ij}$, let $\|\mathbf{A}\|_{1}=\max_{j}\sum_{i}|a_{ij}|$, $\|\mathbf{A}\|_{\infty}=\max_{i}\sum_{j}|a_{ij}|$, and $\left\lVert\mathbf{A}\right\rVert_{\max}=\max_{i,j}|a_{i,j}|$ denote the matrix $1$-norm, $\infty$-norm, and element-wise maximum norm, respectively. Also, let $\|\mathbf{A}\|_{2}$ denote the $\ell_{2}$-operator norm, i.e. the largest singular value of $\mathbf{A}$. For any vector $\mathbf{v}$ with elements $v_{i}$, let $\|\mathbf{v}\|_{2}=\left(\sum_{i}v_{i}^{2}\right)^{1/2}$ and $\|\mathbf{v}\|_{1}=\sum_{i}|v_{i}|$ and $\|\mathbf{v}\|_{\infty}=\max_{i}|v_{i}|$ denote its Euclidean norm, $\ell_{1}$ norm, and $\ell_{\infty}$ norm, respectively. Finally, for an index set $I$, let $\mathbf{v}_{I}$ denote the subvector of $\mathbf{v}$ formed with the elements of $\mathbf{v}$ whose indices $i\in I$, $|I|$ denote its cardinality, and $I^{c}$ denote its complement.
Consider a set of $n$ independent observations $\{(y_{i},\mathbf{X}_{i}^{\top},\mathbf{Z}_{i}^{\top});i=1,\ldots,n\}$ where, without loss of generality, we assume $E(\mathbf{X}_{i})=E(\mathbf{Z}_{i})=\bm{0}$. Each triplet $(y_{i},\mathbf{X}_{i}^{\top},\mathbf{Z}_{i}^{\top})$ follows the multiple index model (1) coupled with the linear instrumental variable model. That is, the full model is given by
@@
where $\mathbf{B}\in\mathbb{R}^{p\times d}$, $\bm{\Gamma}\in\mathbb{R}^{q\times p}$, $\varepsilon_{i}\in\mathbb{R}$ denotes the random noise for the outcome model, and $\mathbf{E}_{i}\in\mathbb{R}^{p}$ denotes the random noise for the instrumental variable model. The vector $(\mathbf{U}_{i}^{\top},\varepsilon_{i})^{\top}\in\mathbb{R}^{p+1}$ is assumed to have mean zero and finite covariance matrix $\bm{\Sigma}$, whose diagonal elements are denoted by $\sigma_{j}^{2}$ for $j=1,\ldots,p+1$. To characterize endogeneity between $\mathbf{X}$ and $\varepsilon$, the off-diagonal elements in the last row and column of $\bm{\Sigma}$ are generally non-zero, while the instrumental variable $\mathbf{Z}_{i}$ satisfies $\mathbf{Z}_{i}\perp(\mathbf{U}_{i}^{\top},\varepsilon_{i})^{\top}$ for all $i=1,\ldots,n$.
We seek to estimate the column space of $\mathbf{B}$, when the dimensions $p$ and $q$ are growing exponentially with the sample size $n$. A common assumption in such settings is that sparsity holds. Specifically: i) each covariate $X_{j}$ depends upon only a few instrumental variables, such that in the instrumental variable model the $j$th column of $\bm{\Gamma}$ is a sparse vector with at most $r_{j}<q$ non-zero elements, where $r=\max_{j=1,\ldots,p}r_{j}$ is bounded, and ii) the multiple index model is sparse such that $y$ depends only on $s<p$ covariates i.e., at most $s$ rows of $\mathbf{B}$ are non-zero row vectors. Equivalently, if $\mathcal{P}(\mathbf{B})=\mathbf{B}(\mathbf{B}^{\top}\mathbf{B})^{-1}\mathbf{B}^{\top}$ denotes the corresponding projection matrix, then only $s$ diagonal elements of $\mathcal{P}(\mathbf{B})$ are non-zero. Without loss of generality, we assume $S=\{1,\ldots,s\}$ i.e., the first $s$ covariates indexes this set of non-zero elements, or, equivalently, the support set of the column space of $\mathbf{B}$. We let $\mathcal{X},\mathcal{Z}$ and $\mathbf{y}$ denote the $n\times p$ matrix of covariates, $n\times q$ matrix of instrumental variables, and $n\times 1$ vector of responses, whose $i$th rows are given by $\mathbf{X}_{i}^{\top}$, $\mathbf{Z}_{i}^{\top}$ and $y_{i}$, respectively. Let $\mathbf{x}_{j}$ and $\mathbf{z}_{j}$ denote the $j$th column of $\mathcal{X}$ and $\mathcal{Z}$, respectively. For the remainder of this article, we use $y,\mathbf{X}$, and $\mathbf{Z}$ to denote the random variables from which $(y_{i},\mathbf{X}_{i}^{\top},\mathbf{Z}_{i}^{\top})$ is drawn.
2.1 Review of sliced inverse regression and Lasso SIR
The sliced inverse regression or SIR estimator (Li,, 1991) of the column space of $\mathbf{B}$ was developed under the assumption that the random noise $\varepsilon$ is independent of the covariates $\mathbf{X}$. The key assumption of SIR is the following linearity condition, which is satisfied when the covariates $\mathbf{X}$ follow an elliptical distribution.
Condition 2.1.
@@ 2.1 Review of sliced inverse regression and Lasso SIR
Under this condition, the population foundation for SIR is that the inverse regression curve $E(\mathbf{X}\mid y)$ is contained in the linear subspace spanned by $\bm{\Sigma}_{\mathbf{X}}\mathbf{B}$ (Theorem 3.1, Li,, 1991). As a result, letting $\{\bm{\eta}^{*}_{1},\ldots,\bm{\eta}^{*}_{d}\}$ denote the set of eigenvectors associated with the $d$ largest eigenvalues of $\bm{\Lambda}^{*}=\operatorname{\text{Cov}}\left\{E(\mathbf{X}\mid y)\right\}$, we have $\bm{\Sigma}_{\mathbf{X}}\bm{\beta}_{k}\propto\bm{\eta}^{*}_{k}$ for $k=1,\ldots,d$. Since $\bm{\Lambda}^{*}=\bm{\Sigma}_{\mathbf{X}}-E\left\{\operatorname{\text{Cov}}(\mathbf{X}\mid y)\right\}$, we can estimate $\bm{\Lambda}^{*}$ from estimates of $\bm{\Sigma}_{\mathbf{X}}$ and $\mathbf{T}=E\left\{\operatorname{\text{Cov}}(\mathbf{X}\mid y)\right\}$. This approach to estimation for SIR is recommended by Li, (1991), and used by many authors in the literature (Tan et al.,, 2018; Zou,, 2006), although it is not the only approach possible (Lin et al.,, 2019).
@@
In more detail, suppose we estimate $\bm{\Sigma}_{\mathbf{X}}$ by the sample covariance $\widehat{\bm{\Sigma}}_{\mathbf{X}}=n^{-1}\mathcal{X}^{\top}\mathcal{X}$. When $n<p$, this sample covariance matrix is not invertible, but the Lasso SIR estimator generally does not require such invertibility as we will see in a moment. Then, partitioning the data into $H$ non-overlapping slices $S_{1},\ldots,S_{h}$ based on the order statistics of $y$, we compute the sample covariance matrices within each slice and average them to form
@@
$\widehat{\mathbf{T}}=\dfrac{1}{H}\sum_{h=1}^{H}\left\{\frac{1}{c-1}\sum_{i\in S_{h}}(\mathbf{X}_{i}-\bar{\mathbf{X}}_{h})(\mathbf{X}_{i}-\bar{\mathbf{X}}_{h})^{\top}\right\},\text{where}\;\bar{\mathbf{X}}_{h}=\dfrac{1}{c}\sum_{i\in S_{h}}\mathbf{X}_{i},$
@@
where $c$ denotes the slice size and we have assumed $n=cH$ to simplify the notation. Zhu et al., (2006) showed that $\widehat{\mathbf{T}}$ can be rewritten in the form
$\widehat{\mathbf{T}}=\dfrac{1}{H(c-1)}\mathcal{X}^{\top}(\mathbf{I}_{H}\otimes\mathbf{P}_{c})\mathcal{X},\text{where}\;\mathbf{P}_{c}=\mathbf{I}_{c}-\dfrac{1}{c}\mathbf{e}_{c}\mathbf{e}_{c}^{\top},$

### v2

@@ Abstract
Sliced inverse regression is a popular sufficient dimension reduction method that identifies a small number of linear transformations of the covariates without losing regression information with the response. In high-dimensional settings, it can be combined with sparsity penalties to achieve simultaneous sufficient dimension reduction and variable selection. Both classical and sparse SIR estimators assume the covariates are exogenous. However, endogeneity can arise in a variety of situations, such as when variables are omitted or are measured with error. In this article, we show that such situations invalidate SIR and lead to inconsistent estimation of the true central subspace. To address this challenge, we propose a two-stage lasso SIR estimator, which first constructs a sparse high-dimensional instrumental variables model to obtain fitted values of the covariates spanned by the instruments, and then applies SIR augmented with a lasso penalty to these fitted values. We establish theoretical bounds for the estimation and selection consistency of the true dimension reduction subspace for the proposed two-stage lasso SIR estimator, allowing the number of covariates and instruments to grow exponentially with the sample size. Simulation studies and applications to two real-world datasets in nutrition and genetics illustrate the superior empirical performance of the proposed estimator, compared with existing methods which ignore endogeneity and/or nonlinearity in the outcome model.
@@ 1 Introduction
Sufficient dimension reduction (Ma and Zhu,, 2013; Li,, 2018, SDR) refers to a class of methods that assume the outcome $y\in\mathbb{R}$ depends on a set of random covariates $\mathbf{X}\in\mathbb{R}^{p}$ only through a small number of linear combinations $\bm{\beta}_{1}^{\top}\mathbf{X},\ldots,\bm{\beta}_{d}^{\top}\mathbf{X}$, with $d\ll p$. These linear combinations are known as sufficient predictors, and retain the full regression information between the response and covariates. More specifically, letting $\mathbf{B}=[\bm{\beta}_{1},\ldots,\bm{\beta}_{d}]$ denote a $p\times d$ matrix, we can formulate SDR as a multiple index model
@@
where $f$ denotes an unknown link function, and $\varepsilon\in\mathbb{R}$ is random noise. The target of estimation is then the column space of $\mathbf{B}$, also known as the central subspace. Since the pioneering work of Li, (1991), a vast literature has developed on different sufficient dimension reduction approaches, ranging from inverse-moment-based and regression-based methods (e.g., Cook and Forzani,, 2008; Bura et al.,, 2022), to forward regression (e.g., Xia et al.,, 2002; Dai et al.,, 2024) and semiparametric techniques (e.g., Zhou and Zhu,, 2016; Sheng and Yin,, 2016). Among the various SDR methods, SIR (Li,, 1991) remains the most widely used due to its simplicity and computational efficiency, and as such continues to be the subject of much research (see Tan et al.,, 2018; Xu et al.,, 2022, among many others). When the number of covariates $p$ is large, it is often (further) assumed each linear combination depends only on a few covariates, i.e., the vectors $\bm{\beta}_{1},\ldots,\bm{\beta}_{d}$ are sparse. Recent literature has thus sought to combine SIR (and other SDR methods) with variable selection to achieve sparse sufficient dimension reduction (e.g., Lin et al.,, 2018, 2019; Nghiem et al., 2023c, ).
@@
Much of the literature dedicated to SIR and many other SDR methods, including those mentioned above, has focused on the setting where all covariates are exogenous, i.e. in model (1), $\mathbf{X}$ and $\varepsilon$ are independent, implying $y\perp\mathbf{X}\mid\mathbf{B}^{\top}\mathbf{X}$, where $\perp$’ denotes statistical independence. However, this assumption may not be reasonable in settings where the random noise $\varepsilon$ is dependent on $\mathbf{X}$, as in the case where the covariates are endogenous. Little has been done to tackle sufficient dimension reduction with endogeneity, with the following two common scenarios. First, omitted-variable bias occurs when one or more (truly important) covariates is not included in the model. In Section 7.1 for instance, we consider a genomic study relating a biological outcome (e.g., body weight) to gene expressions of mice. Outside of gene expressions, there can be potentially unmeasured confounding factors, such as those related to environmental conditions, which influence both gene expression levels and the outcome. Second, covariate measurement error arises when one or more of the covariates are not measured accurately. For instance, in a second data application in Section 7.2, we consider a nutritional study relating an individual’s health outcomes to their long–term diet, collected through 24–hour recall interviews that are well-known to be subject to measurement errors (Carroll et al.,, 2006). Let $\mathbf{X}$ denote the long-term dietary covariates. The observed data can then be modeled as $\mathbf{W}=\mathbf{X}+\mathbf{E}$, with $\mathbf{E}$ denoting an additive measurement error independent of $\mathbf{X}$. In this situation, even when in the true model (1) it holds that $\varepsilon\perp\mathbf{X}$, a naive analysis which replaces $\mathbf{X}$ by $\mathbf{W}$ results in endogeneity, since $y=f(\mathbf{B}^{\top}\mathbf{W}+\mathbf{B}^{\top}\mathbf{E},\varepsilon)$ and $\mathbf{E}$ is correlated with $\mathbf{W}$. This causes naive estimators that ignore measurement errors to be inconsistent.
@@
In this article, we show under general conditions that endogeneity invalidates SIR, leading to inconsistent estimates of the true column space of $\mathbf{B}$. To address this challenge, we propose a two-stage lasso SIR method to estimate the central space in the presence of endogeneity, when valid instrumental variables are available. In the first stage, we construct a sparse multivariate linear model regressing all the covariates against the instruments, and obtain fitted values from this model. In the second stage, we apply a sparse SIR between the response and the fitted values from the first stage to perform simultaneous SDR and variable selection.
@@
Addressing endogeneity via instrumental variables has been extensively studied in the econometrics and statistics literature; see Angrist and Krueger, (2001) for a review. However, classical instrumental variable estimators are mostly studied with a linear model for the outcome, i.e., $y=\mathbf{X}^{\top}\bm{\beta}_{0}+\varepsilon$, and usually compute a two-stage least squares or a limited information maximum likelihood estimator. Hansen et al., (2008) showed in this setting that these estimators are consistent for $\bm{\beta}_{0}$ when the number of instruments $q$ increases slowly with the sample size. More recently, a number of methods have combined instrumental variables with regularization techniques on the covariates and/or instrumental variables to improve their performance in high-dimensional settings e.g., the lasso-type generalized methods of moments (Caner,, 2009; Caner et al.,, 2018), focused generalized methods of moments (Fan and Liao,, 2014), and two-stage penalized least squares (Lin et al.,, 2015); the last of these methods inspires this article. Another line of research has focused on the use of sparsity-inducing penalties to estimate optimal instruments from a large set of a-priori valid instruments or some variation thereof (e.g., Belloni et al.,, 2012; Chernozhukov et al.,, 2015; Windmeijer et al.,, 2019). To the best of our knowledge though, none of these methods allow for the presence of non-linear link functions in the model for the response e.g., the multiple index model in (1). While there is some related literature on single index models with endogeneity (Hu et al.,, 2015; Gao et al.,, 2020), their approach requires approximation of the link function and is not easy to generalize to multiple index models.
@@
Under a setting where both the number of covariates $p$ and the number of instrumental variables $q$ can grow exponentially with sample size $n$, we derive theoretical bounds on the estimation and selection consistency of our proposed two-stage lasso SIR estimator. Simulation studies demonstrate the proposed estimator exhibits superior finite sample performance compared with a single-stage lasso SIR estimator when endogeneity is present but ignored, and the two-stage regularized estimator of Lin et al., (2015) which accounts for endogeneity but assumes the outcome model is linear in nature. We apply the proposed method to two aforementioned real data applications studying the effects of different nutrients on the cholesterol level (covariate measurement error), and for identifying genes that are relevant to the body weight of mice (omitted variable bias).
We establish some notation to be used throughout the paper. For any matrix $\mathbf{A}$ with elements $a_{ij}$, let $\|\mathbf{A}\|_{1}=\max_{j}\sum_{i}|a_{ij}|$, $\|\mathbf{A}\|_{\infty}=\max_{i}\sum_{j}|a_{ij}|$, and $\lVert\mathbf{A}\rVert_{\max}=\max_{i,j}|a_{i,j}|$ denote the matrix $1$-norm, $\infty$-norm, and element-wise maximum norm, respectively. Also, let $\|\mathbf{A}\|_{2}$ denote the $\ell_{2}$-operator norm, i.e. the largest singular value of $\mathbf{A}$. For a square matrix $\mathbf{A}$, we use $\lambda_{j}(\mathbf{A})$ to denote the $j$th largest eigenvalue of $\mathbf{A}$, whose corresponding eigenvector is given by $\bm{\eta}_{j}(\mathbf{A})$. When $j=1$, we suppress the subscript and write $\lambda(\mathbf{A})$ and $\bm{\eta}(\mathbf{A})$. For any vector $\mathbf{v}$ with elements $v_{i}$, let $\|\mathbf{v}\|_{2}=\left(\sum_{i}v_{i}^{2}\right)^{1/2}$, $\|\mathbf{v}\|_{1}=\sum_{i}|v_{i}|$, and $\|\mathbf{v}\|_{\infty}=\max_{i}|v_{i}|$ denote its Euclidean norm, $\ell_{1}$ norm, and $\ell_{\infty}$ norm, respectively. Finally, for an index set $I$, let $\mathbf{v}_{I}$ denote the subvector of $\mathbf{v}$ formed with the elements of $\mathbf{v}$ whose indices $i\in I$, $|I|$ denote its cardinality, and $I^{c}$ denote its complement.
@@ 2 SIR and endogeneity
2.1 Review of SIR and lasso SIR
Consider a set of $n$ independent observations $\{(y_{i},\mathbf{X}_{i}^{\top},\mathbf{Z}_{i}^{\top});i=1,\ldots,n\}$ where, without loss of generality, we assume $E(\mathbf{X}_{i})=E(\mathbf{Z}_{i})={0}$. Each triplet $(y_{i},\mathbf{X}_{i}^{\top},\mathbf{Z}_{i}^{\top})$ follows the multiple index model (1) coupled with a linear instrumental variable model. That is, the full model is given by
@@ 2.1 Review of SIR and lasso SIR
where $\mathbf{B}\in\mathbb{R}^{p\times d}$, $\bm{\Gamma}\in\mathbb{R}^{q\times p}$, $\varepsilon_{i}\in\mathbb{R}$ denotes the random noise for the outcome model, and $\mathbf{U}_{i}\in\mathbb{R}^{p}$ denotes the random noise for the instrumental variable model. The vector $(\mathbf{U}_{i}^{\top},\varepsilon_{i})^{\top}\in\mathbb{R}^{p+1}$ is assumed to have mean zero and a finite covariance matrix $\bm{\Sigma}$ whose diagonal elements are denoted by $\sigma_{j}^{2}$ for $j=1,\ldots,p+1$. To capture endogeneity between $\mathbf{X}$ and $\varepsilon$, the off-diagonal elements in the last row and column of $\bm{\Sigma}$ are generally non-zero. The instrumental variable $\mathbf{Z}_{i}$ satisfies $\mathbf{Z}_{i}\perp(\mathbf{U}_{i}^{\top},\varepsilon_{i})^{\top}$ for all $i=1,\ldots,n$.
We aim to estimate the column space of $\mathbf{B}$ when the dimensions $p$ and $q$ grow exponentially with the sample size $n$. A common assumption in such settings is that sparsity holds. Specifically: (i) each covariate $X_{j}$ depends upon only a few instrumental variables, such that in the instrumental variable model the $j$th column of $\bm{\Gamma}$ is a sparse vector with at most $r_{j}<q$ non-zero elements, where $r=\max_{j=1,\ldots,p}r_{j}$ is bounded, and (ii) the multiple index model is sparse, such that $y$ depends only on $s<p$ covariates i.e., at most $s$ rows of $\mathbf{B}$ are non-zero row vectors. Equivalently, if $\mathcal{P}(\mathbf{B})=\mathbf{B}(\mathbf{B}^{\top}\mathbf{B})^{-1}\mathbf{B}^{\top}$ denotes the corresponding projection matrix, then only $s$ diagonal elements of $\mathcal{P}(\mathbf{B})$ are non-zero. Without loss of generality, we assume $S=\{1,\ldots,s\}$ i.e., the first $s$ covariates indexes this set of non-zero elements, or equivalently the support set of the column space of $\mathbf{B}$.
Let $\mathcal{X}$ denote the $n\times p$ matrix of covariates, whose $i$th rows is $\mathbf{X}_{i}^{\top}$. Similarly, let $\mathcal{Z}$ denote the $n\times q$ matrix of instrumental variables whose $i$th row is $\mathbf{Z}_{i}^{\top}$. Also, let $\mathbf{x}_{j}$ and $\mathbf{z}_{j}$ denote the $j$th column of $\mathcal{X}$ and $\mathcal{Z}$, respectively. For the remainder of this article, we use $y,\mathbf{X}$, and $\mathbf{Z}$ to denote the random variables from which $(y_{i},\mathbf{X}_{i}^{\top},\mathbf{Z}_{i}^{\top})$ is drawn.
The SIR estimator (Li,, 1991) of the column space of $\mathbf{B}$ was developed under the assumption that the random noise $\varepsilon$ is independent of the covariates $\mathbf{X}$. Furthermore, a key prerequisite of SIR is the following linearity condition, which is satisfied when the covariates $\mathbf{X}$ follow an elliptical distribution.
Condition 1.
@@
Under this condition, the population foundation for SIR is that the inverse regression curve $E(\mathbf{X}\mid y)$ is contained in the linear subspace spanned by $\bm{\Sigma}_{\mathbf{X}}\mathbf{B}$ (Theorem 3.1, Li,, 1991). Consequently, if we let $\{\bm{\eta}^{*}_{1},\ldots,\bm{\eta}^{*}_{d}\}=\{\bm{\eta}_{1}(\bm{\Lambda}^{*}),\ldots,\bm{\eta}_{d}(\bm{\Lambda}^{*})\}$ denote the set of eigenvectors associated with the $d$ largest eigenvalues of $\Lambda^{*}=\operatorname{\text{cov}}\left\{E(\mathbf{X}\mid y)\right\}$, then $\bm{\Sigma}_{\mathbf{X}}\bm{\beta}_{k}\propto\bm{\eta}^{*}_{k}$ for $k=1,\ldots,d$. Since $\bm{\Lambda}^{*}=\bm{\Sigma}_{\mathbf{X}}-E\left\{\operatorname{\text{cov}}(\mathbf{X}\mid y)\right\}$, we can estimate $\bm{\Lambda}^{*}$ from some estimates of $\bm{\Sigma}_{\mathbf{X}}$ and $T=E\left\{\operatorname{\text{cov}}(\mathbf{X}\mid y)\right\}$: this approach to SIR is recommended by Li, (1991), and used by many authors in the literature (Tan et al.,, 2018; Zou,, 2006), although we acknowledge it is not the only approach possible (Lin et al.,, 2019).
@@
In more detail, suppose we estimate $\bm{\Sigma}_{\mathbf{X}}$ by the sample covariance matrix $\widehat{\bm{\Sigma}}_{\mathbf{X}}=n^{-1}\mathcal{X}^{\top}\mathcal{X}$. When $n<p$, this sample covariance matrix is not invertible, but critically the lasso SIR estimator defined below does not require such invertibility, as we will see shortly. Next, we partition the data into $H$ non-overlapping slices $S_{1},\ldots,S_{H}$ based on the order statistics of $y$, and compute the sample covariance matrices within each slice and average them as follows
@@
$\widehat{{T}}=\dfrac{1}{H}\sum_{h=1}^{H}\left\{\frac{1}{c-1}\sum_{i\in S_{h}}(\mathbf{X}_{i}-\bar{\mathbf{X}}_{h})(\mathbf{X}_{i}-\bar{\mathbf{X}}_{h})^{\top}\right\},\text{where}\;\bar{\mathbf{X}}_{h}=\dfrac{1}{c}\sum_{i\in S_{h}}\mathbf{X}_{i},$
@@
where $c$ denotes the slice size and we have assumed $n=cH$ to simplify the notation. Zhu et al., (2006) showed that $\widehat{{T}}$ can be rewritten in the form
$\widehat{{T}}=\dfrac{1}{H(c-1)}\mathcal{X}^{\top}(\mathbf{I}_{H}\otimes\mathbf{P}_{c})\mathcal{X},\text{where}\;\mathbf{P}_{c}=\mathbf{I}_{c}-\dfrac{1}{c}{\bm{e}}_{c}{\bm{e}}_{c}^{\top},$